# MDPE: A Multimodal Deception Dataset with Personality and Emotional Characteristics

### v1

@@ Abstract
Deception detection has garnered increasing attention in recent years due to the significant growth of digital media and heightened ethical and security concerns. It has been extensively studied using multimodal methods, including video, audio, and text. In addition, individual differences in deception production and detection are believed to play a crucial role.Although some studies have utilized individual information such as personality traits to enhance the performance of deception detection, current systems remain limited, partly due to a lack of sufficient datasets for evaluating performance. To address this issue, we introduce a multimodal deception dataset MDPE  [^1] The dataset is available at https://github.com/cai-cong/MDPE, licensed under CC BY-NC-SA-4.0 license.. Besides deception features, this dataset also includes individual differences information in personality and emotional expression characteristics. It can explore the impact of individual differences on deception behavior. It comprises over 104 hours of deception and emotional videos from 193 subjects. Furthermore, we conducted numerous experiments to provide valuable insights for future deception detection research. MDPE not only supports deception detection, but also provides conditions for tasks such as personality recognition and emotion recognition, and can even study the relationships between them. We believe that MDPE will become a valuable resource for promoting research in the field of affective computing.
1 Introduction
@@
Generally, deception refers to the act of misleading, tricking, or deceiving others [1]. It involves hiding the truth or presenting false information to create an impression that is not accurate. Deception can take many forms, including both verbal and nonverbal information [2]. And it also occurs in various contexts, such as interpersonal relationships, business, politics, and entertainment. Deception is often considered unethical and can have serious consequences for trust and relationships.
@@
As deception has expanded to other fields such as social media, interviews, online transactions, and deception in daily life,the need arises for a reliable and efficient system to aid the task of detecting deceptive behavior. Many machine learning approaches have been proposed in order to improve the reliability of deception detection systems [3]. In particular, physiological, psychological, visual, linguistic, acoustic, and thermal modalities have been analyzed in order to detect discriminative features and clues to identify deceptive behavior [4, 5, 6, 7]. Video-based deception detection is a current priority in deception research, because behavioral cues can be extracted from videos in a cheaper, faster, and non-invasive manner [8], which is preferable to invasive approaches that extract clues through devices attached to human bodies (e.g., polygraphs). Visual clues of deception include facial emotions, expression intensity, hands and body movements, and microexpressions. These features were shown to be capable of discriminating between deceptive and truthful behavior [9, 10]. Acoustic features took into account the pitch and speaking rate among other measurements to specify whether certain features are associated with an act of deceit [5]. Linguistic features were usually extracted from the language, words usage, and consistency of the statements made by a person [11]. Recently, multimodal analysis has gained a lot of attention due to their superior performance compared to the use of unimodal modalities. In the deception detection field, several multimodal approaches [12, 13, 14, 15] have been suggested to improve deception detection by integrating features from different modalities. This integration created a more reliable system that is not susceptible to factors affecting sole modalities and polygraph tests.
In addition, it is firmly believed that there are individual differences in deception production and detection [16, 17, 18]. Specifically, it includes cognitive level, personality traits, psychological characteristics, and emotional expression. Everyone has different personalities and psychological characteristics, and the expression of emotions is also various. It has been demonstrated through several studies that personality factors and emotional cues play a significant role in subjects’ ability to deceive and detect deception [16, 19]. Emotion is a fundamental aspect of human communication that interacts with cognition, guiding social behavior in both human-to-human interactions and human-computer interactions [20, 21]. Emotional characteristics are important, because deceptive behavior can trigger emotional states, leading to behavioral changes that serve as deceptive clues [9, 22]. However, it is difficult to directly improve the accuracy of deception detection using emotional features [23]. One of the reasons is that emotional expression is also a part of deception. It is usually difficult to detect whether a deceiver’s emotional expression is genuine or disguised.
@@
To address this issue, we propose a multimodal deception dataset MDPE. It not only collects subjects’ deception information, but also personality information and emotional expression information. Each subject was required to conduct another emotional experiment in addition to engaging in deception, in order to obtain their true emotional expression. Although our research was conducted in the laboratory to provide clear and comparable conversations, we provided subjects with effective monetary incentives to detect and generate effective deceptive behavior [16]. To our knowledge, this is the largest multimodal deception dataset in the released dataset and the only deception detection dataset with personality and emotional characteristics.
@@
2 Related Work
Deception Dataset   Pérez-Rosas et al. [12] introduced a new multi-modal deception dataset Real-life Trial having real-life videos of courtroom trials. They demonstrated the use of features from different modalities and the importance of each modality in detecting deception. They also evaluated the performance of humans in deception detection and compared it with their machine learning models. The Box-of-Lies dataset [24] was released with video and audio from a game show, and presents preliminary findings using linguistic, dialog, and visual features. Multiple modalities have been introduced in the hope of enabling more robust detection. Pérez-Rosas et al. [25] introduced a dataset for deception including video and thermal imaging, as well as physiological and audio recordings. Gupta et al. [26] proposed Bag-of-Lies, a multimodal dataset with gaze data for detecting deception in casual settings. Speth Jeremy et al.[27] proposed a multimodal deception database DDPM contains almost 13 hours of recordings of 70 subjects, as well as physiological signals such as thermal video frames and pulse oximeter data. Most studies on deception detection are designed and evaluated on private datasets, typically with relatively small sample sizes, and MDPE dataset addresses these drawbacks. Table 1 compares the sample size and length for existing datasets and MDPE.
Table 1: Comparison of the subject count and length for several databases for deception detection
@@ 2 Related Work
Multimodal Deception Detection   Decades of research in psychology, and deception detection have documented verbal and nonverbal behavioral cues indicative of deceptive communication. Visual cues such as the frequency and duration of eye blinks [28, 29, 30], dilation of pupils [31, 32], and facial muscle movements [33, 34] have been found to distinguish between deceptive and truthful behavior. Vocal cues can be indicative of deception, with deceptive speakers tending to speak with higher and more varied pitch [1, 35], shorter utterances, and less fluency [36, 37] than truthful speakers. Deception also correlates with verbal attributes of speech, with deceivers tending to communicate with less cognitive complexity, fewer self-references, and more words indicative of negative emotions [38, 6]. Mohamed Abouelenien et al. [39] explored a multimodal deception detection approach and integrates multiple physiological, linguistic, and thermal features. They used a decision tree model, to gain insights into the features that are most effective in detecting deceit. Leena Mathur et al. [15] analyzed the discriminative power of features from visual, vocal, and verbal modalities affect for deception detection. They experimented with unimodal Support Vector Machines (SVM) and SVM-based multimodal fusion methods to identify effective features for detecting deception.
@@
Individual Difference Deception   Some studies confirm that some of the five NEO-FFI (Neuroticism-Extraversion-Openness Five-Factor Inventory) dimensions are related to deception [40, 41]. Sarah Ita Levitan et al. [16] reported the role of personality factors derived from the NEO-FFI and of gender, ethnicity and confidence ratings on subjects’ ability to deceive and to detect deception. Justyna Sarzyńska et al. [42] reports correlations between the ability to lie and extraversion, as well as conscientiousness. Personality characteristics are a promising set of information for deception detection, and similarly, emotional characteristics are also important. Joseph P. Gaspar et al. [43] integrate prior theory and research on emotions, emotional intelligence, and deception and introduce a theoretical model. This model explores the interplay between emotional intelligence (the ability to perceive emotions, use emotions, understand emotions, and regulate emotions; and deception. Mircea Zloteanu et al hold strong beliefs about the role of emotional cues in detecting deception, and explored how decoders’ emotion recognition ability and senders’ emotions influence veracity judgements [44]. Joseph P. Gaspar et al. [19] believe that emotions are both an antecedent and a consequence of deception, and they introduce the emotion deception model to represent these relationships. This model broadens their understanding of deception in negotiations and accounts for the important role of emotions in the deception decision process. To our knowledge, MDPE is the only deception detection dataset with personality and emotional characteristics.
3 Dataset
3.1 Materials
@@
We collect our deception dataset using: a sports camera Gopro Hero9 with a resolution of 1920x1080 and a frame rate of 60 fps. The voices of the subjects are also recorded by the built-in microphone of the camera. A Thinkpad laptop was provided to subjects for watching emotion induction videos during the emotion experiment. The experimental place is in a professional recording stdio, and during the data collection process, only the subject and interviewer stay in the room. Some materials were prepared by the Data Collection Coordinator (DCC). The experimental setup is shown in Figure 1.
Figure 1: The example of the place setup for data acquisition.
3.2 Participants
There were 193 subjects in this study, of which 130 were female and 63 were male. They are all native Chinese speakers from different backgrounds. Their age ranged from 18 to 69 years old, and they had various professions including students, workers, teachers, retirees, etc.
Firstly, we segmented the raw video, resulting in 1808 minutes of deceptive video and 4401 minutes of emotional video, totaling 6209 minutes. Each of the 193 subjects provided 24 responses, 9 of which were deceptive. The length of deception videos ranged from 4 minutes to 27 minutes. Each subject had 16 emotional videos, with lengths ranging from 19 minutes to 38 minutes (including the time spent watching emotional induction videos).
Before the experiment began, the subjects were informed of all experimental procedures. The subjects explicitly consented to record their conversation and publish the video data in a scientific conference or journal. And we do not publish any privacy sensitive data, and the anonymity of participants will be guaranteed. All data were collected under a protocol approved by the authors’institution’s Human Subjects Institutional Review Board.
3.3 Procedure
Personality Characteristics Collection: Subjects were required to fill out a Big Five personality questionnaire [45], which consists of 60 questions. Each question was marked with a score indicating whether the descriptions match their own, with 1 indicating strong disagreement and 5 indicating strong agreement. Details can be found in the appendix A.
Emotional Experiment: Subjects were asked to watch 16 emotional induction videos, including two induction videos for each emotion of sadness, happiness, relaxation, surprise, fear, disgust, anger, and neutral. There are a total of 39 induction videos, of which 17 are from the Chinese Emotional Video System (CEVS) [46]. Each video segment has been labeled and evaluated to ensure that it can induce corresponding emotions. Another 22 are from our online collection. Because the CEVS only include six emotions: sadness, happiness, fear, disgust, anger, and neutral, and some videos of the CEVS are outdated and cannot successfully induce corresponding emotions in our pre-experiments. Each video we collected online was annotated by 12 data annotators based on CEVS selection criteria and evaluation methods, and the results showed that each video triggered strong emotions.
@@ 3.3 Procedure
Before the emotional experiment began, the DCC randomly selected 16 induction videos (ensure two videos for each emotion) for the subjects to watch. After watching each video, subjects were required to describe their feelings and then fill out an “Emotional Scale", which quantified 8 emotions. Subjects rated their 8 emotions, ranging from 1 to 5, with 1 indicating no such emotion and 5 indicating the strongest emotion. Details of the emotion scale can be found in the appendix B.
Deception Experiment: The deception data collection process follows DDPM [27]. The interviewer conducted an interview with the subject, asking a total of 24 questions. These 24 questions were jointly formulated by 5 psychology researchers with over 5 years of experience. Before the interview, details of the emotion scale can be found in the appendix C. The DDC randomly selected 9 questions that must lie and hand them over to the subject (the interviewer does not know which 9 questions). The first 3 questions will not be selected, which means that the first three “warm up” questions were always to be answered honestly. They allowed the subject to get settled, and gave the interviewer an idea of the subject’s demeanor when answering a question honestly.
@@
The subject have a maximum of 15 minutes to prepare, and during the preparation process, they must remember these 9 questions and think about how to deceive in the upcoming interview process. During the interview process, when asked these 9 questions, the subject must lie, and when asked the remaining 15 questions, they must tell the truth. Subjects were motivated to deceive successfully through two levels of bonus compensation: if they were able to deceive the interviewer in five or six of the nine deceptive responses, they were given a 150 percent of a base incentive payment; the base payment was doubled if they were successfully deceptive in seven or more questions. In order to collect more indistinguishable deception answers, we encourage subjects to incorporate some truth into lies when answering these deceptive questions.
@@
During the interview process, the interviewer asked 24 questions in random order, and provide their judgment of truthful or deceptive answers to each question. And the interviewer filled out the "Interviewer Judgment Scale", which record the trust level the interviewer thinks of each answer. The trust level was divided into 1-5 points, where 1 represents definitely true and 5 represents definitely false. After the interview, the subject also filled out the "Subject Lie Confidence Scale" and be asked to rate the answer they just lied to. The same score is 1-5, where 1 represents that I have definitely deceived successfully and 5 represents that I have definitely not deceived successfully.
Each subject was required to complete the above three experiments, so that we can obtain their personality , emotional and deceptive characteristics.
4 Benchmark
4.1 Data Preprocessing

### v2

@@ Abstract.
Deception detection has garnered increasing attention in recent years due to the significant growth of digital media and heightened ethical and security concerns. It has been extensively studied using multimodal methods, including video, audio, and text. In addition, individual differences in deception production and detection are believed to play a crucial role. Although some studies have utilized individual information such as personality traits to enhance the performance of deception detection, current systems remain limited, partly due to a lack of sufficient datasets for evaluating performance. To address this issue, we introduce a multimodal deception dataset MDPE  [^1] The dataset is available at https://github.com/cai-cong/MDPE.. Besides deception features, this dataset also includes individual differences information in personality and emotional expression characteristics. It can explore the impact of individual differences on deception behavior. It comprises over 104 hours of deception and emotional videos from 193 subjects. Furthermore, we conducted numerous experiments to provide valuable insights for future deception detection research. MDPE not only supports deception detection, but also provides conditions for tasks such as personality recognition and emotion recognition, and can even study the relationships between them. We believe that MDPE will become a valuable resource for promoting research in the field of affective computing.
multimodal dataset, affective computing, deception detection, personality, emotion
ccs: Human-centered computing HCI design and evaluation methodsccs: Computing methodologies Multi-task learning
1. Introduction
@@
Generally, deception refers to the act of misleading, tricking, or deceiving others (DePaulo et al., 2003). It involves hiding the truth or presenting false information to create an impression that is not accurate. Deception can take many forms, including both verbal and nonverbal information (Burgoon et al., 2021). And it also occurs in various contexts, such as interpersonal relationships, business, politics, and entertainment. Deception is often considered unethical and can have serious consequences for trust and relationships.
@@
As deception has expanded to other fields such as social media, interviews, online transactions, the need arises for a reliable and efficient system to aid the task of detecting deceptive behavior. Many machine learning approaches have been proposed in order to improve the reliability of deception detection systems (Granhag and Hartwig, 2008). In particular, physiological, psychological, visual, linguistic, acoustic, and thermal modalities have been analyzed in order to detect discriminative features and clues to identify deceptive behavior (Feng et al., 2012; Hirschberg et al., 2005; Newman et al., 2003; Rajoub and Zwiggelaar, 2014). Video-based deception detection is a current priority in deception research, because behavioral cues can be extracted from videos in a cheaper, faster, and non-invasive manner (Burzo et al., 2018), which is preferable to invasive approaches that extract clues through devices attached to human bodies (e.g., polygraphs). Visual clues of deception include facial emotions, expression intensity, hands and body movements, and microexpressions. These features were shown to be capable of discriminating between deceptive and truthful behavior (Ekman, 2009; Owayjan et al., 2012). Recently, multimodal analysis has gained a lot of attention due to their superior performance compared to the use of unimodal modalities. In the deception detection field, several multimodal approaches (Pérez-Rosas et al., 2015; Krishnamurthy et al., 2018; Şen et al., 2020; Mathur and Matarić, 2020) have been suggested to improve deception detection by integrating features from different modalities. This integration created a more reliable system that is not susceptible to factors affecting sole modalities and polygraph tests.
Substantial empirical evidence indicates significant individual differences in both deception production and detection capabilities (Levitan et al., 2015; Majumder et al., 2017; Ren et al., 2021). These differences encompass cognitive processing, personality traits, psychological characteristics, and emotional expressivity. Empirical studies confirm that personality factors and emotional cues critically influence subjects’ capacity to deceive and detect deception (Levitan et al., 2015; Gaspar and Schweitzer, 2013). Emotion—a fundamental dimension of human communication—interacts with cognition to guide social behavior across both interpersonal and human-computer interactions (Gordon et al., 2016; Marchi et al., 2015). This relationship is particularly relevant to deception, as deceptive acts can elicit distinctive emotional states that manifest as behavioral clues (Ekman, 2009; Vrij, 2008). However, leveraging emotional features to improve deception detection accuracy remains challenging (Hartwig and Bond Jr, 2014). A primary complicating factor is that emotional expression itself constitutes a core component of deception, making it difficult to discern whether a deceiver’s displayed emotions are genuine or strategically fabricated.
@@
To address this issue, we propose a multimodal deception dataset MDPE. It not only collects subjects’ deception information, but also personality information and emotional information. Each subject was required to conduct another emotional experiment in addition to engaging in deception, in order to obtain their true emotional expression. Although our research was conducted in the laboratory to provide clear and comparable conversations, we provided subjects with effective monetary incentives to detect and generate effective deceptive behavior (Levitan et al., 2015). To our knowledge, this is the largest multimodal deception dataset in the released dataset and the only deception detection dataset with personality and emotional characteristics.
@@
Table 1. Comparison of the subject count and length for several databases for deception detection
@@
2. Related Work
Deception Dataset   Pérez-Rosas et al. (Pérez-Rosas et al., 2015) introduced a new multi-modal deception dataset Real-life Trial having real-life videos of courtroom trials. They demonstrated the use of features from different modalities and the importance of each modality in detecting deception. The Box-of-Lies dataset (Soldner et al., 2019) was released with video and audio from a game show, and presents preliminary findings using linguistic, dialog, and visual features. Multiple modalities have been introduced in the hope of enabling more robust detection. Pérez-Rosas et al. (Pérez-Rosas et al., 2014) introduced a dataset for deception including video and thermal imaging, as well as physiological and audio recordings. Gupta et al. (Gupta et al., 2019) proposed Bag-of-Lies, a multimodal dataset with gaze data for detecting deception in casual settings. Speth Jeremy et al.(Speth et al., 2021) proposed a multimodal deception database DDPM contains almost 13 hours of recordings of 70 subjects, as well as physiological signals such as thermal video frames and pulse oximeter data. Most studies on deception detection are designed and evaluated on private datasets, typically with relatively small sample sizes, and MDPE dataset addresses these drawbacks. Table 1 compares the sample size and length for existing datasets and MDPE.
@@ 2. Related Work
Multimodal Deception Detection   Decades of research in psychology, and deception detection have documented verbal and nonverbal behavioral cues indicative of deceptive communication. Visual cues such as the frequency and duration of eye blinks (Bhaskaran et al., 2011; Fukuda, 2001; Minkov et al., 2012), dilation of pupils (Dionisio et al., 2001; Lubow and Fein, 1996), and facial muscle movements (Hurley and Frank, 2011; Porter et al., 2011) have been found to distinguish between deceptive and truthful behavior. Vocal cues can be indicative of deception, with deceptive speakers tending to speak with higher and more varied pitch (DePaulo et al., 2003; Zuckerman et al., 1981), shorter utterances, and less fluency (Rockwell et al., 1997; Sporer and Schwandt, 2006) than truthful speakers. Deception also correlates with verbal attributes of speech, with deceivers tending to communicate with less cognitive complexity, fewer self-references, and more words indicative of negative emotions (Zhou et al., 2004; Newman et al., 2003). Mohamed et al. (Abouelenien et al., 2016) explored a multimodal deception detection approach and integrates multiple physiological, linguistic, and thermal features. They used a decision tree model, to gain insights into the features that are most effective in detecting deceit. Leena Mathur et al. (Mathur and Matarić, 2020) analyzed the discriminative power of features from visual, vocal, and verbal modalities affect for deception detection. They experimented with unimodal Support Vector Machines (SVM) and SVM-based multimodal fusion methods to identify effective features for detecting deception.
@@
Individual Difference Deception   Some studies confirm that some of the five NEO-FFI (Neuroticism-Extraversion-Openness Five-Factor Inventory) dimensions are related to deception (Ramanaiah et al., 1994; Jakobwitz and Egan, 2006). Sarah Ita Levitan et al. (Levitan et al., 2015) reported the role of personality factors derived from the NEO-FFI and of gender, ethnicity and confidence ratings on subjects’ ability to deceive and to detect deception. Justyna Sarzyńska et al. (Sarzyńska et al., 2017) reports correlations between the ability to lie and extraversion, as well as conscientiousness. Personality characteristics are a promising set of information for deception detection, and similarly, emotional characteristics are also important. Joseph P. Gaspar et al. (Gaspar et al., 2022) integrate prior theory and research on emotions, emotional intelligence, and deception and introduce a theoretical model. This model explores the interplay between emotional intelligence (the ability to perceive emotions, use emotions, understand emotions, and regulate emotions; and deception. Mircea Zloteanu et al hold strong beliefs about the role of emotional cues in detecting deception, and explored how decoders’ emotion recognition ability and senders’ emotions influence veracity judgements (Zloteanu et al., 2021). Joseph P. Gaspar et al. (Gaspar and Schweitzer, 2013) believe that emotions are both an antecedent and a consequence of deception, and they introduce the emotion deception model to represent these relationships. This model broadens their understanding of deception in negotiations and accounts for the important role of emotions in the deception decision process. To our knowledge, MDPE is the only deception detection dataset with personality and emotional characteristics.
Figure 1. Statistical and analytical results of MDPE.
3. Dataset
3.1. Experimental Setup
@@
Equipment: The dataset was collected using a GoPro Hero9 sports camera configured to record video at a resolution of 1920×1080 pixels and a frame rate of 60 frames per second (fps). Audio data were synchronously captured via the camera’s built-in microphone. During the emotional experiment, subjects were provided with a ThinkPad laptop to watch emotion-induction stimuli videos. Data collection took place in a controlled professional recording studio to minimize environmental interference. Only the participant and the interviewer remained in the room during the recordings.
Emotion-Induction Videos: During the emotional induction experiment, subjects were shown a series of emotion-inducing videos designed to evoke specific emotional states, including sadness, happiness, relaxation, surprise, fear, disgust, anger, and neutrality. A total of 39 videos were utilized, with 17 collected from the Chinese Emotional Video System (CEVS) (Pengfei Xu and Luo, 2010). Each video segment in the CEVS has been professionally labeled and evaluated to ensure its effectiveness in eliciting the corresponding emotional responses. However, the CEVS only includes six emotions: sadness, happiness, fear, disgust, anger, and neutrality, excluding relaxation and surprise. Additionally, some of the CEVS videos failed to reliably evoke the intended emotional responses during our pre-experiment. This may be attributed to shifts in aesthetic preferences over time, resulting in reduced emotional resonance among contemporary viewers. To address this problem, an additional 22 videos were collected from online sources. These videos were annotated by 12 independent data annotators according to the same criteria and annotation methods as those used in the CEVS. The results showed that each video successfully elicited strong emotional responses.
Deception Questions: During the deception experiment, participants were asked a set of 24 ”deception questions”. These questions were developed by a panel of five psychology researchers, each with over five years of experience, drawing upon theoretical frameworks such as the Fraud Triangle Theory and Rational Choice Theory. To ensure comprehensiveness, the questions were designed to integrate interdisciplinary perspectives from psychology, criminology, and sociology, thereby capturing diverse dimensions of deceptive behavior. Some questions were specifically designed to reflect emerging trends in deceptive practices, addressing aspects potentially overlooked by conventional methodologies. The initial question set underwent a preliminary round of testing, during which participant feedback was collected and incorporated into subsequent revisions. The finalized version of the deception questions is provided in Appendix C.
Personnel: The study involved two distinct personnel roles: the interviewer and the Data Collection Coordinator (DCC). Interviewers were researchers with at least three years of experience in psychology and had received specialized training in deception detection techniques. Their responsibilities included conducting interviews, making real-time judgments regarding participants’ truthfulness, and completing the Interviewer Judgment Scale. The data collection coordinator will assist with the execution of the study, including the preparation of materials and other related tasks.
Other Material: The Interviewer Judgment Scale assessed the interviewer’s perception of each subject’s response credibility, rated on a 5-point Likert scale (1 = definitely true, 5 = definitely false). After the interview, subjects completed the Confidence Scale for Lying to self-assess their deceptive performance, also using a 5-point Likert scale (1 = I definitely deceived successfully, 5 = I definitely did not deceive successfully). During the emotional experiment, subjects rated their experience of eight specific emotions on an Emotional Scale, using a 5-point scale (1 = no such emotion, 5 = strongest emotion). Details can be found in Appendix B.
3.2. Procedure
Each subject was required to complete the following three experiments: personality, emotion and deception experiment.
Personality Characteristics Collection: Subjects were required to complete a Big Five personality questionnaire (Zhang et al., 2022), which consists of 60 items. Each item was rated on a Likert scale from 1 to 5, with 1 indicating strong disagreement and 5 indicating strong agreement, reflecting the degree to which each statement matched the subject’s own characteristics. Based on the scoring methodology, the Big Five personality traits—openness, conscientiousness, extraversion, agreeableness, and neuroticism—were assessed for each subject. The full questionnaire and scoring methodology are provided in Appendix A.
@@ 3.2. Procedure
Emotional Experiment: The DCC randomly selected 16 induction videos (ensure 2 videos for each emotion) for the subjects to watch. After watching each video, subjects were required to describe their feelings and then fill out an Emotional Scale.
@@
Deception Experiment: The deception data collection process follows DDPM (Speth et al., 2021). The DCC randomly selected 9 questions that must lie and hand them over to the subject (the interviewer does not know which 9 questions). The first 3 questions will not be selected, which means that the first three “warm up” questions were always to be answered honestly. They allowed the subject to get settled, and gave the interviewer an idea of the subject’s demeanor when answering a question honestly. The subject have a maximum of 15 minutes to prepare, and during the preparation process, they must remember these 9 questions and think about how to deceive in the upcoming interview process. During the interview process, when asked these 9 questions, the subject must lie, and when asked the remaining 15 questions, they must tell the truth. Subjects were motivated to deceive successfully through two levels of bonus compensation: if they were able to deceive the interviewer in five or six of the nine deceptive responses, they were given a 150% of a base incentive payment; the base payment was doubled if they were successfully deceptive in seven or more questions. In order to collect more indistinguishable deception answers, we encourage subjects to incorporate some truth into lies when answering these deceptive questions.
@@
During the interview process, the interviewer asked 24 questions in random order, and provide their judgment of truthful or deceptive answers to each question. And the interviewer filled out the Interviewer Judgment Scale. And after the interview, the subject also filled out the Subject Lie Confidence Scale.
3.3. Statistic and Analysis
A total of 193 subjects took part in this study, comprising 130 females and 63 males, all of whom were native Chinese speakers. Their ages ranged from 18 to 69 years, and their occupations included students, laborers, teachers, retirees, and others. Each subject contributed responses to 24 deception questions, yielding a total of 1737 deceptive and 2895 truthful responses. In addition, each participant provided 16 emotion-inducing videos, resulting in a total of 3088 emotional video recordings. Following data collection, the raw video recordings were segmented. The duration of individual deceptive video clips ranged from 4 to 27 minutes, while emotional videos ranged from 19 to 38 minutes, which included the time spent watching emotion-induction materials. In total, 1808 minutes of deceptive video and 4401 minutes of emotional video were obtained, amounting to 6209 minutes of video content.
A preliminary analysis of the data reveals the distribution of successful deception attempts across all subjects (Figure 1 (a)). The majority of participants had success rates in the 3-6 times range. Figure 1 (b) displays the average number of successful deception by personality category. Subjects were categorized into high and low groups for each Big Five personality trait based on mean scores. Analysis indicates no significant difference in deception success rates between high/low Neuroticism or Openness. However, significantly higher deception success rates were observed among individuals with low Extraversion, high Agreeableness, and low Conscientiousness. The finding regarding low Extraversion may be attributed to a tendency towards greater caution and reduced likelihood of revealing vulnerabilities through exaggeration. Individuals scoring high in Agreeableness, characterized by cooperativeness and trustfulness, may more easily gain trust and goodwill, potentially lowering interviewer vigilance. Similarly, those with low Conscientiousness typically place less emphasis on rules, obligations, and social norms. These patterns align with established perspectives from disciplines such as self-control theory and general crime theory (Collins and Schmidt, 1993; Detert et al., 2008; Gottfredson and Hirschi, 1990). Figure 1 (c) presents the average intensity of each emotion expressed during the emotion experiment, comparing subjects with high versus low deception success rates. Individuals exhibiting higher deception success rates demonstrated heightened emotional expressivity. This may reflect a greater capacity among successful deceivers to understand and simulate others’ emotions, resulting in a more sensitive and reactive emotional system. This observation is consistent with existing theories linking emotion and cognition (Austin et al., 2007; Grandey, 2003).
3.4. Ethics Review and License
Before the experiment began, the subjects were informed of all experimental procedures. The subjects explicitly consented to record their conversation and publish the video data in a scientific conference or journal. And we do not publish any privacy-sensitive data, and the anonymity of participants will be guaranteed. All data were collected under a protocol approved by the authors’institution’s Human Subjects Institutional Review Board.
Additionally, we restrict the use of this dataset under the license of CC BY-NC-SA-4.0, requiring researchers to use our dataset responsibly. And commercial usage is prohibited.
4. Benchmark
4.1. Data Preprocessing